OpenStudy (anonymous):

Integral of Sec^3x-secx dx This is the last part of a much bigger problem (;

OpenStudy (anonymous):

Integral (sec^3(x) dx) First, split sec^3(x) into sec(x) and sec^2(x). Integral (sec(x) sec^2(x) dx) Now, use integration by parts. Let u = sec(x). dv = sec^2(x) dx du = sec(x)tan(x). v = tan(x) Integral (sec^3(x) dx) = sec(x)tan(x) - Integral (sec(x)tan^2(x) dx) Use the identity tan^2(x) = sec^2(x) - 1. Integral (sec^3(x) dx) = sec(x)tan(x) - Integral (sec(x)[sec^2(x) - 1] dx) Distribute the sec(x). Integral (sec^3(x) dx) = sec(x)tan(x) - Integral( (sec^3(x) - sec(x)) dx) Now, separate into two integrals. Integral (sec^3(x) dx) = sec(x)tan(x) - [Integral (sec^3(x) dx) - Integral (sec(x) dx)] Distribute the minus over the brackets. Integral (sec^3(x) dx) = sec(x)tan(x) - Integral (sec^3(x) dx) + Integral (sec(x) dx) Here's the part which gets tricky; we're going to move - Integral (sec^3(x) dx) to the left hand side of our equation, resulting in TWO of them. 2 Integral (sec^3(x) dx) = sec(x)tan(x) + Integral (sec(x) dx) And we know what the integral of sec(x) is (we stated it above). 2 Integral (sec^3(x) dx) = sec(x)tan(x) + ln|sec(x) + tan(x)| All we have to do now is divide everything by 2, which is the same as multiplying everything by (1/2). Integral (sec^3(x) dx) = (1/2)sec(x)tan(x) + (1/2)ln|sec(x) + tan(x)| And don't forget to add the constant. = (1/2)sec(x)tan(x) + (1/2)ln|sec(x) + tan(x)| + C

OpenStudy (anonymous):

Im so sorry one last thing. It was supposed to be 9 times that integral. So on each piece I pull out of the integral will it also be multiplied by the constant 9? For example would it break into 9uv-int u'v?